Jefferson Lab HALL A Logbook for November 07 - All EXCEPT Automatic Entries
Main INDEX
Friday
217306 11/30/07 22:59 Boeglin Compton, BPM, BCM strip chart
217305 11/30/07 22:57 Boeglin Ozone Strip chart
217304 11/30/07 22:57 Adaq Hall A Tools
217297 11/30/07 22:40 mcnulty Target screens
217292 11/30/07 22:31 boeglin Beam sport size OK
217286 11/30/07 22:21 boeglin beam spot size
217280 11/30/07 21:32 boeglin Spot size
217277 11/30/07 21:20 boeglin beam spot
217272 11/30/07 19:00 Boeglin Raster Snapshot
217269 11/30/07 17:12 Jack Checking the status of the LQ2 power supply
217268 11/30/07 17:09 Jack (Techs) Left Q2 FuG power supply status
217267 11/30/07 16:46 LeRose Setting Left Q2
217266 11/30/07 16:09 LeRose Shift Summary
217264 11/30/07 14:26 A. Camsonne Increased NaI HV by 100V
217263 11/30/07 14:25 Y. Zhang target figure
217262 11/30/07 14:24 LeRose Rad/Ozone/BCM
217261 11/30/07 11:31 Yoomin ps4 in run 3069, 3070
217260 11/30/07 10:13 Y. Zhang Left Cerenkov Calibration
217259 11/30/07 08:31 LeRose Controlled Access
217254 11/30/07 07:53 R. Michaels owl shift summary
217240 11/30/07 06:41 R. Michaels tools
217235 11/30/07 06:20 R. Michaels P0 change
217230 11/30/07 05:55 R. Michaels P0 changes
217221 11/30/07 05:10 R. Michaels reboot adaql3
217214 11/30/07 04:51 R. Michaels tools again (R-arm closer)
217208 11/30/07 04:33 R. Michaels tools (starting R-arm runs)
217202 11/30/07 03:57 R. Michaels Ozone/Rad Mon
217201 11/30/07 03:56 R. Michaels BPM/BCM
217200 11/30/07 03:55 R. Michaels Compton/OTR
217197 11/30/07 03:34 R. Michaels empty pion rej. histos
217194 11/30/07 03:22 R. Michaels P0 change
217193 11/30/07 03:13 H Bandara Target Status
217172 11/30/07 01:51 R. Michaels beam returns
217167 11/30/07 00:32 R. Michaels polarity change / beam update
217165 11/30/07 00:01 Boeglin Shift summary
Thursday
217164 11/29/07 23:18 M. Stevens Tech/Left dipole polarity change
217159 11/29/07 22:12 mcnulty Moved target from He --> dummy
217155 11/29/07 21:59 mcnulty Moved target from Pb --> He
217153 11/29/07 21:55 V. Sulkosky Re: Problem with physics analysis for run 3032
217150 11/29/07 21:43 Boeglin BCM BPM Charts
217149 11/29/07 21:42 Boeglin Ozone Radiation Strip Chart
217148 11/29/07 21:41 Boeglin Hall A Tools Screen
217147 11/29/07 21:30 mcnulty Moved target from H2 --> Pb
217140 11/29/07 21:13 mcnulty Moved target from C --> H2
217137 11/29/07 21:08 Boeglin Problem with physics analysis for run3032
217134 11/29/07 20:49 mcnulty Moved target from Fe --> C
217131 11/29/07 20:44 mcnulty Target screens
217124 11/29/07 20:02 mcnulty Moved target from C --> Fe
217119 11/29/07 19:36 mcnulty Moved target from dummy --> C
217114 11/29/07 19:28 mcnulty Moved target from H2 --> dummy
217097 11/29/07 18:02 mcnulty Moved target from C --> Fe
217090 11/29/07 17:35 mcnulty Moved target from H2 --> dummy
217083 11/29/07 17:17 mcnulty Moved target from Pb --> H2
217078 11/29/07 16:47 mcnulty Moved target from He --> Pb
217069 11/29/07 16:17 mcnulty Moved target from Pb --> He
217066 11/29/07 16:10 Holmstrom Shift summary
217046 11/29/07 14:40 Holmstrom Momentum Change Left
217040 11/29/07 14:19 Y. Zhang BCM/BPM page
217039 11/29/07 14:18 Y. Zhang hall a tools page
217038 11/29/07 14:17 Y. Zhang target page
217028 11/29/07 13:13 A. Camsonne NaI channel 60
217027 11/29/07 13:12 Holmstrom Momentum Change
217022 11/29/07 12:57 A. Camsonne Pedestal subtraction off for NaI channel 61 ( histo 60 )
217008 11/29/07 11:54 Y. Zhang NaI bad channel
216966 11/29/07 08:05 jpchen change left p to 217.441 MeV
216965 11/29/07 08:04 jpchen owl shift summary
216952 11/29/07 07:22 jpchen tools screen
216951 11/29/07 07:22 jpchen Radiation/Ozone
216950 11/29/07 07:21 jpchen BPM/BCM
216949 11/29/07 07:20 jpchen Compton/OTR
216939 11/29/07 06:28 jpchen change left to 233.214 MeV, right to 268.275 MeV
216916 11/29/07 04:51 jpchen target screens
216910 11/29/07 04:26 jpchen change left to 250.131 MeV
216895 11/29/07 03:12 jpchen radiation/ozone
216894 11/29/07 03:12 jpchen BPM/BCM
216893 11/29/07 03:11 jpchen Compton/OTR strip charts
216892 11/29/07 03:10 jpchen General tools screen
216890 11/29/07 03:03 jpchen change left to 268.275 MeV, right to 287.735 MeV
216866 11/29/07 01:34 jpchen change left to 287.735 MeV
216848 11/29/07 00:19 jpchen change both to 308.607MeV
216841 11/29/07 00:02 boeglin Shift summary
Wednesday
216837 11/28/07 23:55 mcnulty Moved target from H2 --> Pb
216832 11/28/07 23:45 mcnulty Moved target from dummy --> H2
216827 11/28/07 23:36 mcnulty Moved target from C --> dummy
216824 11/28/07 23:27 Boeglin BPM and BCM Strip charts
216823 11/28/07 23:26 Boeglin Ozone and Radiation Strip Chart
216822 11/28/07 23:24 Boeglin Hall A Tools
216819 11/28/07 23:08 mcnulty Moved target from Fe --> C
216816 11/28/07 22:56 Boeglin Righ Arm Dipole not quite stable during run 21224
216805 11/28/07 22:12 mcnulty Moved target from C --> Fe
216784 11/28/07 20:39 mcnulty Moved target from dummy --> Pb
216780 11/28/07 20:24 mcnulty Target Screens
216764 11/28/07 19:25 mcnulty Moved target from Pb --> dummy
216763 11/28/07 19:22 V. Sulkosky Pion Rejector layer 1, ADC 21
216754 11/28/07 18:55 mcnulty Moved target from Fe --> He
216743 11/28/07 18:10 mcnulty Moved target from Pb --> dummy
216726 11/28/07 17:44 mcnulty Moved target from He --> Pb
216717 11/28/07 17:14 mcnulty Moved target from C --> Fe
216708 11/28/07 16:54 mcnulty Moved target from H2 --> dummy
216703 11/28/07 16:42 mcnulty Moved target fr0m Pb --> H2
216698 11/28/07 16:31 mcnulty Moved target from He --> Pb
216693 11/28/07 16:10 mcnulty Moved target from Fe --> He
216687 11/28/07 15:57 Nilanga shift summary
216682 11/28/07 15:41 L. Zhu raster ok now
216669 11/28/07 15:05 L. Zhu raster not good
216662 11/28/07 13:30 L. Zhu beam is back
216661 11/28/07 13:14 Ole Hansen How to check quota
216660 11/28/07 12:21 L. Zhu target ready
216659 11/28/07 10:04 V. Sulkosky Deadtime RHRS runs 21175 to 21193
216658 11/28/07 09:36 R. Michaels adaq quota doubled (thanks CC)
216653 11/28/07 08:43 V. Sulkosky LHRS Pion Rejector HV
216651 11/28/07 07:59 BT, Hu shift summary
216650 11/28/07 05:52 Y. Zhang radiation/ozone strip chart
216649 11/28/07 05:47 Y.Zhang BCM/BPM graph
216648 11/28/07 05:45 Y.Zhang tools page
216647 11/28/07 05:15 jpchen CHL crashed, targets warmed up, He loop vented.
216646 11/28/07 05:12 Y. Zhang CHL crash
216645 11/28/07 05:08 Y. Zhang CHL crashed
216644 11/28/07 05:06 Y.Zhang CHL crashed
216643 11/28/07 05:01 BT, Hu BPM and BCM chart
216642 11/28/07 04:58 BT, Hu compton rate and OTR chart
216641 11/28/07 04:27 Y.Zhang physics replay
216638 11/28/07 03:56 BT, Hu momentum change
216627 11/28/07 03:26 A. Camsonne Low statistics run but need to chekc since pedestal run was done a bit earlier
216616 11/28/07 02:59 BT, Hu change left p to 540.394MeV/c
216607 11/28/07 02:26 Y. Zhang replay
216602 11/28/07 02:03 BT, Hu momentum change
216585 11/28/07 01:00 R. Michaels quota command
216580 11/28/07 00:48 jpchen Helium target vented and re-cooled at 5-6 pm.
216579 11/28/07 00:42 R. Michaels why halog broke (fixed now, I hope)
216576 11/28/07 00:34 R. Michaels halog test #454241
0 11/28/07 00:25 no_name no_key
216574 11/28/07 00:24 Adaq test2
0 11/28/07 00:22 no_name no_key
0 11/28/07 00:20 no_name no_key
0 11/28/07 00:15 no_name no_key
0 11/28/07 00:15 no_name no_key
0 11/28/07 00:13 no_name no_key
0 11/28/07 00:09 no_name no_key
0 11/28/07 00:08 no_name no_key
0 11/28/07 00:07 no_name no_key
0 11/28/07 00:05 no_name no_key
Tuesday
0 11/27/07 23:23 no_name no_key
0 11/27/07 23:23 no_name no_key
0 11/27/07 23:20 no_name no_key
0 11/27/07 23:19 no_name no_key
0 11/27/07 23:16 no_name no_key
0 11/27/07 23:12 no_name no_key
0 11/27/07 23:11 no_name no_key
0 11/27/07 23:09 no_name no_key
0 11/27/07 23:09 no_name no_key
0 11/27/07 21:46 no_name no_key
0 11/27/07 20:12 no_name no_key
0 11/27/07 20:10 no_name no_key
0 11/27/07 20:06 no_name no_key
216549 11/27/07 19:18 A. Camsonne NaI HV increased
0 11/27/07 18:18 no_name no_key
216543 11/27/07 17:42 A. Camsonne Pedestal subraction on
216531 11/27/07 16:40 A. Camsonne Pedestal subtraction off
216523 11/27/07 16:02 Boeglin Shift Summary
216521 11/27/07 15:54 M. Stevens Left Q2 overtemp problem
216520 11/27/07 14:45 hbandara cryotarg reset
216519 11/27/07 14:26 spiegel Hall A Gas
216518 11/27/07 11:25 hbandara Target status
216514 11/27/07 08:10 BT, Hu shift sumarry
216497 11/27/07 06:08 BT, Hu change right p to 0.89041GeV/c
216484 11/27/07 05:28 BT, Hu change left p to 952.73MeV/c
216477 11/27/07 05:03 Y. Zhang NaI replay
216462 11/27/07 04:14 BT, Hu general tools' GUI
216459 11/27/07 04:07 BT, Hu compton and OTR strip chart
216458 11/27/07 04:07 Y. Zhang target chart
216457 11/27/07 04:06 BT, Hu BMP and bcm chart
216452 11/27/07 03:51 BT, Hu Change left p to 1.09079GeV/c
216427 11/27/07 02:38 BT, Hu change left p to 1.24884GeV/c
216414 11/27/07 01:20 jpchen move left angle to 40 degree
216413 11/27/07 01:20 jpchen change left p to 1.42980 GeV
216404 11/27/07 00:49 jpchen change left p to 0.15 GeV
216401 11/27/07 00:36 jpchen right angle to 40 degree, p to 1.33626 GeV
216399 11/27/07 00:31 jpchen Swing Shift Summary
216397 11/27/07 00:20 jpchen change left p to 0.3 GeV
216391 11/27/07 00:02 jpchen Compton/OTR
Monday
216388 11/26/07 23:53 jpchen right optics study completed
216386 11/26/07 23:43 jpchen change left p to 0.5 GeV
216376 11/26/07 23:15 jpchen change left p to 1 GeV
216365 11/26/07 22:29 jpchen change right p to 1.57132 (-4%)
216362 11/26/07 22:19 jpchen left optics co,mplete, move P to 1.288023 GeV
216353 11/26/07 21:43 jpchen change right p to 1.60406 GeV (-2%)
216352 11/26/07 21:41 jpchen change left p to 1.57132 GeV (-4%)
216342 11/26/07 21:14 jpchen change left p to 1.60406 GeV(-2%)
216339 11/26/07 21:08 jpchen tools screen, right dipole set
216338 11/26/07 21:06 jpchen radiation/ozone strip chart
216337 11/26/07 21:05 jpchen OTR, BPM
216336 11/26/07 21:04 jpchen Compton
216330 11/26/07 20:45 jpchen change both p to 1.63679 GeV (0%)
216317 11/26/07 18:52 jpchen MCC is ready to send beam
216316 11/26/07 18:51 jpchen right dipole is ready
216315 11/26/07 18:27 jpchen beam off to tune for Hall B
216314 11/26/07 18:27 jpchen change both p to 1.66953 GeV (+2%)
216305 11/26/07 17:37 jpchen beam back, right dipole set, raster off.
216304 11/26/07 16:20 Xinhu Yan Target Charts
216302 11/26/07 16:11 jpchen change both p to 1.70226 (4%) for optics study
216301 11/26/07 16:11 A. Camsonne Day shift summary
216292 11/26/07 15:44 A. Camsonne Production finished MCC working on MARC8 box supply
216287 11/26/07 15:33 A. Camsonne Right septum scope
216268 11/26/07 14:40 V. Sulkosky Ozone/Radiation Stripchart Entries
216261 11/26/07 14:12 A. Camsonne Ozone / radiation
216260 11/26/07 14:11 A.Camsonne Compton /radiation strips
216257 11/26/07 13:41 A. Camsonne Started right arm momentum change
216256 11/26/07 13:37 A. Camsonne Left arm started momentum change to 1.201 GeV/c
216225 11/26/07 12:20 A. Camsonne Asked to restore energy
216206 11/26/07 11:40 A. Camsonne Energy shift
216201 11/26/07 11:26 A. Camsonne NaI HV changed
216192 11/26/07 11:00 A. Camsonne Started changed left momentum to 1.24294
216188 11/26/07 10:46 Yoomin Target chart
216183 11/26/07 10:21 Yoomin spot_L
216181 11/26/07 10:16 A. Camsonne Started right arm momentum change to 1.288 GeV/c
216172 11/26/07 09:56 A. Camsonne Left arm set to 1.28802 GeV/c
216171 11/26/07 09:52 A. Camsonne Starting left arm momemtum change to 1288.023 MeV/c
216161 11/26/07 09:29 Yoomin Wrong entries for run 2802-2807 -- momentum setting
216155 11/26/07 09:14 Yoomin Run 2804, 21109 are junk runs
216154 11/26/07 09:11 A. Camsonne Right arm almost settled
216151 11/26/07 09:07 A. Camsonne SLI still does not look good resuming production
216142 11/26/07 08:55 A. Camsonne Moving to empty position to check SLI
216141 11/26/07 08:53 spiegel Gas system
216136 11/26/07 08:37 A. Camsonne Left arm momentum set to 1.3331 GeV/c
216135 11/26/07 08:34 A. Camsonne Starting momentum change left arm
216127 11/26/07 08:07 A. Camsonne Starting right arm momentum change
216114 11/26/07 07:35 Seonho Choi Summary of Runs (OWL Monday 11/26/2007)
216097 11/26/07 06:23 Y. Zhang target chart
216086 11/26/07 05:47 Seonho Choi Radiation/Ozone chart
216085 11/26/07 05:46 Seonho Choi BPM/BCM Stripcharts
216082 11/26/07 05:44 Seonho Choi Hall A Tools page
216081 11/26/07 05:43 Seonho Choi Compton Rates
216068 11/26/07 05:07 Seonho Choi Final raster
216057 11/26/07 04:51 Seonho Choi Raster adjustments
216052 11/26/07 04:31 Seonho Choi Target to Fe beam right
216051 11/26/07 04:28 Seonho Choi beam position adjusted
216050 11/26/07 04:27 Seonho Choi Harp scan (1H03B) after adjustments
216049 11/26/07 04:26 Seonho Choi Harp scan (1H03A) after adjustments
216048 11/26/07 04:17 Seonho Choi HARP Scan just after Compton tuning
216047 11/26/07 04:11 Seonho Choi Tune finished through Compton Chicane
216046 11/26/07 01:20 Seonho Choi Compton rates
216044 11/26/07 00:13 A. Camsonne giving up on SLI starting tune in Compton chicane
Sunday
216043 11/25/07 23:38 A. Camsonne SLI
216042 11/25/07 23:15 A. Camsonne OTR OK !
216041 11/25/07 23:11 A. Camsonne OTR/SLI
216040 11/25/07 23:08 A. Camsonne Beam aspect ratio
216037 11/25/07 22:18 A. Camsonne Beam postion OK
216036 11/25/07 22:15 A. Camsonne OTR not displaying
216035 11/25/07 22:09 A.Camsonne Hall A SLI
216018 11/25/07 21:08 Xinhu Yan Target Charts
216017 11/25/07 21:07 A. Camsonne NaI Off
216014 11/25/07 16:54 A. Camsonne/ NaI cable swapped
216012 11/25/07 15:31 jpchen cryotarget cool-down completed
216011 11/25/07 14:43 V. Sulkosky LHRS and RHRS Momenta Set
216010 11/25/07 13:20 A. Camsonne Arcnet board moved to HV 3
216009 11/25/07 12:31 Folts guards
216007 11/25/07 12:02 jpchen Replaced a N2 gas bottle for NaI detector
216005 11/25/07 11:41 jpchen Cooldown cryo target started
216004 11/25/07 10:41 V. Sulkosky Reboot IOCHA14
216003 11/25/07 10:21 V. Sulkosky Left Q2 Still Up
216001 11/25/07 00:48 Xinhu Yan Left arm Cerenkov Calibration
216000 11/25/07 00:14 Xinhu Yan Right arm Cerenkov Calibration
Saturday
215996 11/24/07 13:59 jpchen Changed N2 gas bottle for NaI
215995 11/24/07 12:50 Xinhu Yan Calibration for Total Shower target=He,angle=120
215994 11/24/07 12:39 Xinhu Yan Calibration for Total Shower
215993 11/24/07 12:39 Xinhu Yan Calibration for Total Shower
Friday
215992 11/23/07 13:36 spiegel Hall A Gas system
Thursday
215990 11/22/07 20:05 A. Camsonne TS11 crashed
215989 11/22/07 20:03 A. Camsonne NaI nitrogen bottle swapped
215988 11/22/07 12:58 Xinhu Yan Update the scaler config file
215987 11/22/07 12:58 Xinhu Yan Update the scaler config file
215986 11/22/07 12:56 Xinhu Yan Update the scaler config file
215985 11/22/07 12:55 Xinhu Yan Update the scaler config file
Wednesday
215982 11/21/07 17:16 V. Sulkosky LHRS and RHRS at 15 degs.
215981 11/21/07 16:46 folts Q2 overtemp
215980 11/21/07 16:35 Xinhu Yan Bcm Scaler Calibration Constants for right arm
215979 11/21/07 15:37 V. Sulkosky BPM positions after Ped Update
215978 11/21/07 13:48 V. Sulkosky BPM Pedestals (Corrected Link)
215977 11/21/07 13:45 V. Sulkosky BPM Pedestals
215974 11/21/07 11:43 V. Sulkosky Re: BMPA Attenuator Module (swap completed)
215973 11/21/07 10:46 V. Sulkosky Alarms in Hall A Counting House
215972 11/21/07 10:19 Xinhu Yan The constants are for left arm scaler
215971 11/21/07 10:19 Xinhu Yan The constants are for left arm scaler
Tuesday
215957 11/20/07 23:37 Xinhu Yan Bcm Scaler Calibration Constants
215956 11/20/07 23:35 Xinhu Yan Bcm Scaler Calibration Constants
215955 11/20/07 23:21 Ole Hansen Re: disk full warning
215954 11/20/07 17:09 V. Sulkosky BPMA Attenuator Change LHRS (Temporarily)
215953 11/20/07 16:25 Xinhu Yan The Coefficient for Bcm has been changed
215952 11/20/07 15:05 paschke PREX: tested beam modulation system
215951 11/20/07 13:29 A. Camsonne NaI status
215950 11/20/07 13:25 A. Camsonne NaI plots
215949 11/20/07 13:24 A. Camsonne NaI plots
215948 11/20/07 10:26 R. Michaels /halla/scratch nearly full
215946 11/20/07 09:04 Higinbotham Short Day Shift
215944 11/20/07 07:19 BT, Hu owl shift summary
215943 11/20/07 07:14 BT, Hu BCM gui
215942 11/20/07 07:13 BT, Hu BMP
215941 11/20/07 07:11 BT, Hu compton gui stopped undating around 6;00
215940 11/20/07 07:09 BT, Hu beam off at 7;00
215937 11/20/07 07:08 BT, Hu change left to 0 and right to 0
215928 11/20/07 06:43 BT, Hu change left to 114.624MeV/c
215897 11/20/07 04:51 BT, Hu change left to 140.025MeV/c
215891 11/20/07 04:09 BT, Hu General tools gui
215890 11/20/07 04:07 BT, Hu current
215889 11/20/07 04:06 BT, Hu compton gui
215887 11/20/07 04:05 BT, Hu BMP
215884 11/20/07 03:47 BT, Hu change right to 154.765MeV/c
215867 11/20/07 02:58 BT, Hu change left to 171.056
215862 11/20/07 02:46 jpchen Checked Nitrogen bottles for NaI
215861 11/20/07 02:44 BT, Hu compton trip
215858 11/20/07 02:33 jpchen Compton is up ad running again
215857 11/20/07 02:31 jpchen Compton DAQ crashed
215840 11/20/07 00:52 BT, Hu left to 247.713 right to 189.062
215828 11/20/07 00:01 Ole Hansen SWING shift summary
Monday
215821 11/19/07 23:45 Ole Hansen Choppy beam
215808 11/19/07 22:13 Xinhu Yan Bcm calibration constants for EPIC
215807 11/19/07 22:11 Xinhu Yan Bcm calibration constants for EPIC
215794 11/19/07 21:15 V. Sulkosky Re: L/R Asymmetry
215793 11/19/07 21:15 Ole Hansen BCM/BPM strip charts
215792 11/19/07 21:14 Ole Hansen Ozone strip chart
215791 11/19/07 21:13 Ole Hansen Hall A Tools screenshot
215785 11/19/07 20:40 Ole Hansen spot++ run 2726
215782 11/19/07 20:28 Ole Hansen HRSL to 284 MeV, HRSR to 265 MeV
215781 11/19/07 20:27 Ole Hansen L/R asymmetry(cont)
215771 11/19/07 20:05 Ole Hansen L/R asymmetry
215766 11/19/07 19:44 Ole Hansen Changed outputRight.odef
215763 11/19/07 19:40 Ole Hansen Patched C++ analyzer
215750 11/19/07 18:44 Norum rates discrepancy
215743 11/19/07 18:16 Ole Hansen Left HRS to 305.6 MeV
215734 11/19/07 17:28 Ole Hansen HRSR Q2/Q3 cycled
215733 11/19/07 17:19 Ole Hansen HRSR to 305.6 MeV
215732 11/19/07 17:18 Ole Hansen Target 10cm dummy
215729 11/19/07 17:11 Ole Hansen Target H2
215726 11/19/07 16:57 Ole Hansen Left Q2/Q3 cycled
215725 11/19/07 16:54 Ole Hansen Hall back to beam permit
215724 11/19/07 16:51 Ole Hansen Left Q2 1400A max
215722 11/19/07 16:25 Ole Hansen Positron runs finished
215711 11/19/07 16:05 LeRose Shift Summary (Day)
215702 11/19/07 15:36 Higinbotham BCM & BPM Strip Charts
215701 11/19/07 15:34 Higinbotham Ozone Chart (presently 0.025 ppm)
215696 11/19/07 15:28 Higinbotham Tools Screen & BCM
215691 11/19/07 15:10 LeRose Spot
215654 11/19/07 12:10 LeRose Hall A Tools kin 18 left
215637 11/19/07 11:10 Saha He added to Positron Runs
215626 11/19/07 10:29 LeRose Hall A Tools
215623 11/19/07 10:02 LeRose End Controlled Access
215622 11/19/07 09:10 LeRose Controlled Access
215605 11/19/07 08:03 jpchen owl shift summary
215597 11/19/07 07:53 jpchen Left P to 327.794 MeV
215586 11/19/07 07:15 jpchen Compton/OTR charts
215583 11/19/07 07:14 jpchen BPM/BCM charts
215582 11/19/07 07:14 jpchen radiation/ozone chart
215581 11/19/07 07:13 jpchen Hall A Tools screen
215568 11/19/07 06:13 jpchen Change Left to 351.572 MeV, Right to 327.794 MeV
215543 11/19/07 04:39 jpchen change left P to 377.074 MeV
215540 11/19/07 04:23 jpchen target charts
215539 11/19/07 04:22 jpchen BPM/BCM charts
215538 11/19/07 04:20 jpchen OTR/Compton charts
215537 11/19/07 04:18 jpchen radiation/ozone charts
215536 11/19/07 04:18 jpchen tools screen: right dipole set
215529 11/19/07 03:28 jpchen Started to change right P to 351.572 MeV
215514 11/19/07 02:40 jpchen Left -> 404.427 MeV
215497 11/19/07 01:06 jpchen Left->433.763 MeV, Right ->377.074 MeV
Sunday
215477 11/18/07 23:50 mcnulty moved target from Fe --> C
215463 11/18/07 21:57 mcnulty moved target from He --> Fe
215461 11/18/07 21:46 A. Camsonne Nitrogen bottle swapped
215458 11/18/07 21:40 mcnulty moved target from Pb --> He
215451 11/18/07 21:10 mcnulty moved target from H2 --> Pb
215450 11/18/07 21:10 mcnulty moved target from dummy --> H2
215441 11/18/07 20:45 mcnulty moved target from C --> dummy
215440 11/18/07 20:44 mcnulty target charts
215437 11/18/07 20:34 mcnulty moved target from Fe --> C
215436 11/18/07 20:33 mcnulty moved target from C --> Fe
215431 11/18/07 20:25 mcnulty R-arm Q1 continually fluctuating by small amount
215424 11/18/07 19:48 mcnulty moved target from dummy --> C
215421 11/18/07 19:30 mcnulty moved target from He --> dummy
215414 11/18/07 19:19 mcnulty moved target from Pb --> He
215413 11/18/07 19:07 mcnulty moved target from Fe --> Pb
215408 11/18/07 18:47 mcnulty moved target from C --> Fe
215401 11/18/07 18:22 mcnulty moved target from dummy --> C
215394 11/18/07 18:07 mcnulty moved target from Pb to dummy
215391 11/18/07 17:51 mcnult moved target from Pb --> H2
215381 11/18/07 17:17 mcnulty moved target from Pb --> He
215378 11/18/07 16:59 mcnulty moved target from dummy --> Pb
215374 11/18/07 16:40 mcnulty moved target from C to dummy
215372 11/18/07 16:39 mcnulty moved target from Fe --> C
215366 11/18/07 16:21 mcnulty moved target from He --> Fe
215362 11/18/07 16:02 X. Zheng day shift summary
215359 11/18/07 15:52 X. Zheng Compton rate stripchart
215338 11/18/07 14:49 K.Allada run #2622 is a good run
215329 11/18/07 14:10 X. Zheng RF problem
215328 11/18/07 14:09 X. Zheng radiation_ozone chart
215327 11/18/07 14:01 X. Zheng target screen #2
215326 11/18/07 14:01 X. Zheng target screen #1
215325 11/18/07 13:50 X. Zheng BCM screen
215324 11/18/07 13:49 X. Zheng Hall A tools screen
215321 11/18/07 13:39 X. Zheng p_right setting was 708.181
215294 11/18/07 11:49 A. Camsonne Nitrogen gas check
215281 11/18/07 11:13 X. Zheng new raster setting
215275 11/18/07 10:56 X. Zheng New beam position and Compton rate
215270 11/18/07 10:13 X. Zheng Compton rate very high
215261 11/18/07 09:42 X. Zheng Rates at the new kinematics (Fe-right target)
215256 11/18/07 09:22 K.Allada target to beam right iron
215255 11/18/07 09:20 X. Zheng raster size on scope
215254 11/18/07 09:18 X. Zheng spot size
215247 11/18/07 09:04 X. Zheng beam position okay
215240 11/18/07 08:38 X. Zheng harp scan result (good)
215239 11/18/07 08:16 Solvignon Shift summary
215237 11/18/07 06:16 Solvignon Tuned beam
215234 11/18/07 05:34 Solvignon Hall A tools for next settings
215233 11/18/07 05:16 Solvignon Left arm at 90deg
215232 11/18/07 04:25 Solvignon Controlled access
215231 11/18/07 04:14 Solvignon Right arm at 90deg. Left arm stuck at 84.5deg
215230 11/18/07 03:37 Yoomin spot_R
215229 11/18/07 03:33 Adaq spot_L
215226 11/18/07 01:58 Solvignon Pass change
215207 11/18/07 00:44 Yoomin Target charts
215193 11/18/07 00:00 mcnulty moved target from Pb --> He
Saturday
215188 11/17/07 23:34 mcnulty moved target from H2 --> Pb
215184 11/17/07 23:28 mcnulty moved target from Fe --> H2
215178 11/17/07 22:49 mcnulty moved target from C --> Fe
215173 11/17/07 22:30 mcnulty moved target from dummy --> C
215168 11/17/07 21:56 mcnulty target charts
215167 11/17/07 21:55 mcnulty moved target from He --> dummy
215164 11/17/07 21:44 mcnulty moved target from Pb --> He
215143 11/17/07 20:10 mcnulty moved from Fe --> C
215132 11/17/07 19:01 mcnulty moved target from C --> Fe
215127 11/17/07 18:39 mcnulty moved target from Fe --> C
215126 11/17/07 18:34 mcnulty moved target from C --> Fe
215125 11/17/07 18:33 mcnulty moved target from dummy --> C
215114 11/17/07 17:53 A. Camsonne Compton trips
215107 11/17/07 17:19 mcnulty moved target from Pb --> dummy
215104 11/17/07 17:09 mcnulty moved target from Helium to Pb
215097 11/17/07 16:46 mcnulty moved target from empty --> Helium
215094 11/17/07 16:29 A. Camsonne Raster size
215089 11/17/07 16:23 A. Camsonne Spot with no raster
215088 11/17/07 16:22 A. Camsonne Beam size check after optics tune
215083 11/17/07 16:14 A. Camsonne Raster
215076 11/17/07 16:09 A. Camsonne Compton chart
215073 11/17/07 16:07 A. Camsonne Compton rates after Hall B tune
215071 11/17/07 15:57 Xinhu Yan Day Shift Summary
215068 11/17/07 15:29 X. Zheng wrong entries (215051/2)
215067 11/17/07 15:29 A. Camsonne Beam taken away for Hall B optics tune
215064 11/17/07 15:21 A. Camsonne Tail in raster
215053 11/17/07 14:55 A. Camsonne Beam size
215052 11/17/07 14:53 X. Zheng Reboot Roc 13 and 14
215051 11/17/07 14:26 X. Zheng Hall back to beam permit now
215050 11/17/07 14:00 A. Camsonne Harp scan results
215049 11/17/07 13:43 A. Camsonne Compton rates after beam resumed
215042 11/17/07 10:43 Xinhu Yan Magnet status
215041 11/17/07 10:40 Xinhu Yan BPM &BCM status
215040 11/17/07 10:39 xinhu Yan Hall A Control Tool
215039 11/17/07 10:38 Xinhu Yan recycle Q2&Q3 for right arm
215038 11/17/07 10:19 Xinhu Yan Target Charts
215022 11/17/07 08:08 V. Sulkosky Owl Shift Summary
215016 11/17/07 07:31 V. Sulkosky Raster Size
215008 11/17/07 07:03 V. Sulkosky Updated Runplan
215007 11/17/07 06:46 fansler RD1 start up fault/liquid level fault
215006 11/17/07 06:41 V. Sulkosky LHRS Q2 Current Limit
215005 11/17/07 06:17 V. Sulkosky RHRS Q1 back up, dipole still Down
215004 11/17/07 06:08 fansler LQ2 OVERTEMP FAULT
215003 11/17/07 05:22 V. Sulkosky Cryo Problem, maybe ESR
215002 11/17/07 05:01 V. Sulkosky LHRS Magnet Stripchart
215001 11/17/07 04:50 V. Sulkosky BCM D3 Rate on RHRS
215000 11/17/07 03:47 R. Michaels re: DAQ repair (belated)
214999 11/17/07 03:24 V. Sulkosky Contolled Access Again
214992 11/17/07 02:55 V. Sulkosky Techs: Left Q2 Overtemperature Fault
214982 11/17/07 02:04 V. Sulkosky Status Update
214973 11/17/07 01:12 Yoomin Target Chart
214970 11/17/07 00:36 atrig Trigger-Download
214964 11/17/07 00:27 atrig Trigger-Download
214963 11/17/07 00:26 atrig Trigger-Download
214960 11/17/07 00:01 Solvignon Shift summary
Friday
214955 11/16/07 23:35 brads Webcam working, bubbler still bubbling
214952 11/16/07 22:46 A. Camsonne ROC 3 power supply -15V flaky
214951 11/16/07 22:27 Solvignon Controlled access
214949 11/16/07 22:06 mcnulty moved target from C --> dummy
214948 11/16/07 22:03 mcnulty Target Charts
214947 11/16/07 22:02 Solvignon BCM/BPM stripcharts
214946 11/16/07 22:02 Solvignon Radiation/ozone stripcharts
214945 11/16/07 22:01 Solvignon Compton and OTR strip charts
214944 11/16/07 21:59 A. Camsonne Dead time issue ROC3 again
214943 11/16/07 21:59 Solvignon Hall A tools
214942 11/16/07 21:51 Solvignon NaI bubbler webcam not working
214937 11/16/07 21:39 mcnulty moved target from dummy --> C
214931 11/16/07 21:26 Solvignon Spot_R
214923 11/16/07 21:05 A. Camsonne power cycled ROC3
214913 11/16/07 20:44 A. Camsonne more on ROC3 dead time
214909 11/16/07 20:26 A. Camsonne TDC ROC3 slot 4 replaced
214906 11/16/07 19:49 A. Camsonne High deadtime issue
214904 11/16/07 19:44 mcnulty moved target from Pb --> dummy
214891 11/16/07 19:09 mcnulty moved target from Hydrogen --> Pb
214885 11/16/07 18:42 mcnulty target moved from Dummy --> Hydrogen
214884 11/16/07 18:42 E.Chudakov Left DAQ problem
214871 11/16/07 18:05 mcnulty Target moved from C --> Dummy
214866 11/16/07 17:56 mcnulty moved from Fe to Carbon target
214865 11/16/07 17:52 Adaq NaI HV channel 136 - Histo 133
214860 11/16/07 17:30 Solvignon Best BPM positions we can get
214859 11/16/07 17:29 E.Chudakov Raster size
214858 11/16/07 17:28 Solvignon Left collimator is now out
214854 11/16/07 16:48 Solvignon Tuned beam
214847 11/16/07 10:55 Higinbotham Right Arm Run Plan Error
214846 11/16/07 09:24 Higinbotham Change to 5th Pass
214845 11/16/07 09:24 jpchen owl + Early day shift summary
214815 11/16/07 07:31 jpchen moved left collimator to 6 msr
214814 11/16/07 07:29 jpchen optics study with left spectrometer
214805 11/16/07 07:06 jpchen left p to 2202.075 MeV
214802 11/16/07 06:56 jpchen Left P to 2248 MeV
214800 11/16/07 06:50 Xinhu Yan Target charts
214798 11/16/07 06:42 jpchen left sieve slits in
214797 11/16/07 06:39 jpchen change right P to 1914.282 MeV, Left to 2293 MeV
214788 11/16/07 06:01 jpchen main tools
214786 11/16/07 06:01 jpchen BPM/BCM
214785 11/16/07 06:00 jpchen radiation/ozone strip chart
214784 11/16/07 05:59 jpchen Compton/OTR strip charts
214764 11/16/07 04:47 Xinhu Yan target screens
214750 11/16/07 03:50 jpchen changed left to P=2053.142 MeV
214747 11/16/07 03:38 jpchen target screens
214746 11/16/07 03:37 jpchen tools screen
214731 11/16/07 02:55 jpchen Dead time -18%?
214727 11/16/07 02:41 jpchen FFB working now
214725 11/16/07 02:35 jpchen Compton/OTR
214724 11/16/07 02:34 jpchen radiation/ozone strip chart
214723 11/16/07 02:33 jpchen BPM/BCM screens
214722 11/16/07 02:32 Xinhu Yan Target He-->>Pb
214719 11/16/07 02:11 jpchen Change both to P=2022.075 MeV
214716 11/16/07 02:03 jpchen T2 much higher than T4
214713 11/16/07 01:57 jpchen FFB and orbit lock
214712 11/16/07 01:55 Xinhu Yan Target Pb-->>He
214709 11/16/07 01:44 jpchen Beam status
214706 11/16/07 01:31 Xinhu Yan Target Dummy-->>Pb
214697 11/16/07 00:56 Xinhu Target Fe-->>C12
214688 11/16/07 00:08 Xinhu Yan Target Charts
Thursday
214684 11/15/07 23:59 Solvignon Shift summary
214681 11/15/07 23:46 Shabestari Target screen
214678 11/15/07 23:18 Solvignon Raster on Iron_left
214675 11/15/07 23:04 Solvignon BCM/BPM stripcharts
214674 11/15/07 23:03 Solvignon Radiation/Ozone stripchart
214673 11/15/07 23:02 Solvignon Compton stripchart
214672 11/15/07 23:01 Solvignon Hall A tools
214671 11/15/07 22:38 Solvignon Raster check
214668 11/15/07 22:32 Shabestari Target: Fe
214652 11/15/07 21:19 A. Camsonne Compton rates
214646 11/15/07 20:29 A. Camsonne Compton rates
214643 11/15/07 20:06 A. Camsonne Left arm checkout while Compton tuning
214636 11/15/07 19:16 A. Camsonne Right arm HV crate back on
214625 11/15/07 16:51 Solvignon Quads cycled
214622 11/15/07 14:13 Higinbotham Right LeCroy Problem
214612 11/15/07 11:41 A. Camsonne NaI HV changed
214611 11/15/07 08:20 Higinbotham Change to 3rd Pass
214610 11/15/07 08:06 jpchen Owl shift summary
214609 11/15/07 08:06 jpchen raster size with Fe
214599 11/15/07 07:49 jpchen Acceptance study
214590 11/15/07 07:27 A. Camsonne Dead time check
214589 11/15/07 07:27 A. Camsonne Dead time check
214570 11/15/07 06:37 jpchen summary of Pb target study with raster size
214559 11/15/07 06:02 jpchen right spectrometer set by 5:47
214555 11/15/07 05:56 A. Camsonne Lead target scan plots
214547 11/15/07 05:36 jpchen Left P set to 339.636 MeV
214545 11/15/07 05:26 jpchen Completed production data at 60 degree.
214533 11/15/07 04:33 jpchen target screens
214523 11/15/07 03:56 jpchen right dipole at 2x10-3 now
214520 11/15/07 03:47 jpchen Left to 305.957 MeV
214514 11/15/07 03:15 jpchen Compton strip chart: DAQ restarted
214512 11/15/07 03:10 jpchen change right p to 276.818 MeV
214505 11/15/07 02:53 jpchen Compton/OTR
214502 11/15/07 02:51 jpchen radiation/ozone
214501 11/15/07 02:50 jpchen BPM/BCM
214500 11/15/07 02:49 jpchen tools screen
214499 11/15/07 02:45 Yoomin Correction for run 2354
214494 11/15/07 02:35 jpchen right dipole set, change left p to 328.151 MeV
214481 11/15/07 02:08 jpchen right dipole not set yet
214479 11/15/07 02:04 jpchen right dipole set
214475 11/15/07 01:43 jpchen started to move right P to 305.957 MeV
214466 11/15/07 01:06 jpchen Left --> 351.955 MeV
214460 11/15/07 00:35 Yoomin dp plot of Fe 120 degree, 0.845 E_beam
Wednesday
214452 11/14/07 23:57 Nilanga Shift Summary
214449 11/14/07 23:49 Shabestari Target: H2
214444 11/14/07 23:35 Shabestari Target : Fe --> 12C --> Dummy
214438 11/14/07 23:13 Shabestari Target: Fe
214432 11/14/07 22:09 HBandara Beam trip
214429 11/14/07 21:47 Shabestari Target: 12C
214426 11/14/07 21:22 Shabestari Target: dummy
214419 11/14/07 20:56 Shabestari Target screen
214418 11/14/07 20:55 Shabestari Target: H2
214413 11/14/07 20:44 Shabestari Target: Pb
214406 11/14/07 20:29 Hbandara Dead time
214405 11/14/07 20:26 Shabestari 4He target
214396 11/14/07 20:08 Hbandara Harp scan
214394 11/14/07 18:27 jpchen right dipole set
214392 11/14/07 18:03 jpchen Magnets all set except right dipole
214391 11/14/07 17:48 jpchen Target cooldown on the way. Setting up magnets
214390 11/14/07 17:35 jpchen Spectrometer magnets cryo flow are stable.
214388 11/14/07 13:58 Saha Hall A Arc Energy Measurement
214387 11/14/07 09:44 Higinbotham HRS Magnets Alive
214385 11/14/07 05:33 Meziani Comment on beam availability
214384 11/14/07 05:24 Meziani Spot_R
214383 11/14/07 05:22 Meziani spot_L
214382 11/14/07 05:17 Meziani Summary of Bull's eye study
214375 11/14/07 04:54 Meziani Bull's eye scan (+2,-2)
214370 11/14/07 04:45 meziani Bull's eye scan (-2,-2)
214369 11/14/07 04:40 meziani Back on Beam permit
214366 11/14/07 03:49 Meziani Comment on Bull's eye (-2,2)
214365 11/14/07 03:44 meziani Controlled access
214364 11/14/07 03:39 meziani Bull's eye scan (+2,+2)
214359 11/14/07 03:10 Meziani Bull's eye (-2,2)
214350 11/14/07 02:25 meziani Bull's eye scan (0,0)
214349 11/14/07 02:07 jpchen Pb target information
214343 11/14/07 01:38 A. Camsonne Compton tune by MCC done
214342 11/14/07 01:26 A. Camsonne Injector tune improved
214341 11/14/07 01:10 A. Camsonne Compton background
Tuesday
214340 11/13/07 20:37 Jack Right dipole work done today
214339 11/13/07 19:30 V. Sulkosky LHRS BPM Part II
214338 11/13/07 19:01 V. Sulkosky LHRS BPM Problematic History
214337 11/13/07 17:30 Sirish Compton Back Online
214335 11/13/07 15:46 Higinbotham Shift Summary (day shift)
214334 11/13/07 15:31 A. Camsonne Pedestal subtraction off
214331 11/13/07 12:54 Higinbotham ESR Crash (Target & Magnets Warm)
214330 11/13/07 11:19 V. Sulkosky Re: Cavity Monitors
214329 11/13/07 10:50 B. Hahn Cavity Monitors
214327 11/13/07 09:01 Higinbotham Beam Studies
214323 11/13/07 08:47 Yoomin Pb Target problem
214311 11/13/07 08:10 Z. Meziani Shift Summary
214303 11/13/07 07:46 Z. Meziani HRSL new momentum
214302 11/13/07 07:35 Yoomin Plots for L-HRS runs with 120 deg, e0=0.845GeV, Target=Carbon right
214283 11/13/07 06:02 Z. Meziani New momentum for both spectrometers
214258 11/13/07 04:35 Z. Meziani BPM/BCM charts
214257 11/13/07 04:34 Z. Meziani Ozone/Radiation chart
214254 11/13/07 04:32 Z. Meziani Hall A tools
214253 11/13/07 04:26 Z. Meziani Momentum change (left arm)
214244 11/13/07 03:40 Z. Meziani Momentum change on right spectrometer
214229 11/13/07 02:37 Z. Meziani TSO disconnect
214219 11/13/07 01:58 jpchen raster x/y for Pb scan (cont)
214216 11/13/07 01:57 jpchen raster x/y for Pb scan (cont)
214215 11/13/07 01:55 jpchen raster x/y for Pb postion scan
214214 11/13/07 01:54 jpchen raster x/y for Pb position scan runs
214211 11/13/07 01:47 jpchen missing plots for 2206 and 2208
214209 11/13/07 01:45 jpchen missing plots
214205 11/13/07 01:26 jpchen Pb position scan y_target plots
214204 11/13/07 01:23 Z. Meziani Momentum change of HRSL and HRSR mometum
214199 11/13/07 01:11 jpchen LH2 cell parameters
214182 11/13/07 00:30 jpchen LH2 low pressure alarm due to solinoid valve closed
214178 11/13/07 00:04 R. Michaels re: ABU script
214177 11/13/07 00:01 Ole Hansen Shift summary
Monday
214164 11/12/07 23:13 Ole Hansen HRSL to 616 MeV
214153 11/12/07 22:21 folts tech/ left SH AC
214150 11/12/07 22:01 Ole Hansen BTA problems fixed
214149 11/12/07 22:00 Ole Hansen HRSR to 574 MeV
214148 11/12/07 21:59 Ole Hansen HRSL to 661 MeV
214135 11/12/07 21:21 Ole Hansen Rad/Ozone screen
214133 11/12/07 21:20 Ole Hansen BPM/BCM strip charts
214132 11/12/07 21:19 Ole Hansen Hall A Tool screen
214128 11/12/07 21:09 Ole Hansen BTA / abuloop restarted
214121 11/12/07 20:50 Ole Hansen HRSR settled at 661 MeV
214120 11/12/07 20:49 Ole Hansen HRSL to 708 MeV
214119 11/12/07 20:49 Ole Hansen BTA update
214114 11/12/07 20:27 Ole Hansen Erroneous start run info
214113 11/12/07 20:21 Ole Hansen BTA software problem
214106 11/12/07 19:45 Ole Hansen HRSR momentum slightly off
214105 11/12/07 19:43 Ole Hansen HRSR to 661 MeV/c
214104 11/12/07 19:42 Ole Hansen HRSL to 760.38 MeV/c
214091 11/12/07 19:01 brads Changes to BPMA hardware
214082 11/12/07 18:39 Ole Hansen spot++ for run 2332
214077 11/12/07 18:20 Ole Hansen HRSL fixed. Ready for beam.
214076 11/12/07 17:58 V. Sulkosky BPMA Attenuator Change LHRS
214075 11/12/07 17:57 V. Sulkosky BPMA Attenuator Change LHRS
214074 11/12/07 17:50 Ole Hansen Right arm NMR woes
214069 11/12/07 16:47 Ole Hansen Left Arm AC not working
214068 11/12/07 16:24 Ole Hansen Left arm detector hut too hot
214067 11/12/07 16:22 Ole Hansen Spectrometers at 60 deg
214065 11/12/07 16:02 Higinbotham Shift Summary
214064 11/12/07 16:01 LeRose Radiation Monitor
214063 11/12/07 14:30 Higinbotham Positron Run Summary
214046 11/12/07 13:44 LeRose Target Screens
214038 11/12/07 13:21 LeRose H2 target adjusted dE/dx
214010 11/12/07 12:03 gomez ArchiverViewer of right dipole momentum change Nov 11, 2007 @ ~ 12:55 PM
214003 11/12/07 11:38 Higinbotham Left Arm Momentum Runs 2308 - 2314
213989 11/12/07 11:04 A. Camsonne Run 2307 Gas Cerenkov with positive polarity
213982 11/12/07 10:50 LeRose H2 fill high pressure alarm change
213968 11/12/07 10:12 Higinbotham Left Arm Quads Cycled
213964 11/12/07 09:40 Higinbotham Left Arm Set to Positive
213962 11/12/07 08:05 Meziani Shift Summary
213939 11/12/07 05:05 meziani Target charts
213936 11/12/07 04:59 meziani Current chart
213935 11/12/07 04:57 meziani Hall A tools
213930 11/12/07 04:38 meziani BPM chart during the beam scan on the PB target
213929 11/12/07 04:35 meziani Ozone/Radiation monitor Chart for the beam scan
213928 11/12/07 04:32 Z. Meziani Beam position scan on Pb
213907 11/12/07 02:24 Z. Meziani update
213892 11/12/07 00:57 jpchen ROC1 and right DAQ reset
213890 11/12/07 00:01 cusanno Swing Shift Summary
213889 11/12/07 00:01 jpchen Loop 3 (LH2) pressure
Sunday
213885 11/11/07 23:39 morgenstern Lead target in
213884 11/11/07 23:36 cusanno new momentum setting poL = 113.37 MeV
213881 11/11/07 23:12 morgenstern Iron target is in
213878 11/11/07 22:45 cusanno H2 pressure ok now
213875 11/11/07 22:39 cusanno Carbon target in
213874 11/11/07 22:38 cusanno Increasing pressure on Loop 3
213872 11/11/07 22:22 garibaldi Gas Pressure Alarm
213868 11/11/07 22:10 cusanno L-H2 target in
213866 11/11/07 22:01 garibaldi screenshot of cryotarget controls
213865 11/11/07 21:59 garibaldi ESR stripchart
213864 11/11/07 21:58 garibaldi target stripcharts
213863 11/11/07 21:55 cusanno OTR and Compton
213862 11/11/07 21:54 cusanno Ozone and Radiation
213861 11/11/07 21:54 cusanno BPM and BCM
213860 11/11/07 21:52 cusanno Hall A Tools
213859 11/11/07 21:41 cusanno p0L = 125.30 MeV, p0R = 138.49 MeV
213858 11/11/07 21:17 cusanno Right Arm P0 regulation OFF, last Right Arm data are junk?
213850 11/11/07 20:34 cusanno Target restored, asked for beam permit
213849 11/11/07 19:33 jpchen cryotarget power tripped and Helium target vented
213848 11/11/07 18:15 cusanno target problem, controlled access
213845 11/11/07 16:45 garibaldi Lead target in
213838 11/11/07 16:29 garibaldi H2 target
213837 11/11/07 16:28 cusanno IOC of Right Arm rebooted
213834 11/11/07 16:12 cusanno IOC Right Spectrometer Lost
213833 11/11/07 16:02 S. Sirca Day shift SUMMARY
213823 11/11/07 14:57 Timing scintillator
213820 11/11/07 14:54 S. Sirca Stabilization time for HRSR dipole
213814 11/11/07 13:34 S. Sirca Right HRS S2 L vs. R PMT TDCs
213792 11/11/07 10:33 S. Sirca BPM/BCM stripcharts
213791 11/11/07 10:20 K.Allada target screens
213782 11/11/07 09:28 S. Sirca Radiation/ozone snapshot
213781 11/11/07 09:27 S. Sirca Hall A General Tools
213778 11/11/07 09:23 K.Allada target screens
213776 11/11/07 09:14 S. Sirca Nicer OTR DataCube snapshot
213773 11/11/07 08:59 S. Sirca OTR DataCube snapshot
213768 11/11/07 08:01 Meziani Shift Summary
213759 11/11/07 07:40 Yoomin Radiation Ozone Strip
213754 11/11/07 07:08 Yoomin Bad beam current
213753 11/11/07 06:57 meziani Beam Current chart
213748 11/11/07 06:33 Yoomin Target chart
213719 11/11/07 03:41 Yoomin Target Chart
213716 11/11/07 03:32 meziani Spot_R
213715 11/11/07 03:30 meziani SPOT_L
213710 11/11/07 03:16 meziani Magnet levels Left
213709 11/11/07 03:15 meziani Magnet levels right
213706 11/11/07 03:14 Meziani BPM/BCM charts
213705 11/11/07 03:12 Meziani Hall A tools+Current chart
213694 11/11/07 02:26 A. Camsonne NaI HV increased
213673 11/11/07 00:04 Solvignon Swing shift summary
Saturday
213671 11/10/07 23:54 Solvignon Empty histogram in left arm pion rejector
213668 11/10/07 23:36 Solvignon Target moved to Pb
213663 11/10/07 23:19 Solvignon Target moved to He
213658 11/10/07 22:32 Solvignon Target moved to Iron
213653 11/10/07 21:43 Solvignon Target moved to carbon
213647 11/10/07 21:33 Solvignon Target moved from LH2 to Dummy
213640 11/10/07 20:25 Solvignon Target moved to Pb
213632 11/10/07 20:00 Solvignon Target moved to He
213626 11/10/07 19:32 Solvignon Target moved to lead
213622 11/10/07 19:20 Solvignon Target moved to LH2
213617 11/10/07 19:12 Solvignon Target screen (II)
213616 11/10/07 19:11 Solvignon Target screen (I)
213615 11/10/07 19:09 Solvignon Target moved to Dummy
213613 11/10/07 19:06 Solvignon BPM/BCM stripcharts
213612 11/10/07 19:05 Solvignon Radiation/ozone stripchart
213610 11/10/07 19:04 Solvignon Compton stripchart
213609 11/10/07 19:02 Solvignon Hall A tools
213608 11/10/07 18:53 Solvignon spot++
213602 11/10/07 18:39 Solvignon Target moved to carbon
213595 11/10/07 18:15 A. Camsonne Pedestal subtraction removed on NaI channel 141
213593 11/10/07 18:00 Solvignon Target moved to Iron
213591 11/10/07 17:44 Solvignon NaI detector channel 141
213589 11/10/07 17:34 Solvignon Target moved to carbon
213586 11/10/07 17:24 Solvignon Target moved to dummy
213583 11/10/07 17:12 Solvignon Target moved to LH2
213578 11/10/07 16:41 Solvignon Target moved to Pb
213565 11/10/07 16:17 Solvignon Target moved to He
213562 11/10/07 16:02 S. Sirca Day shift SUMMARY
213514 11/10/07 11:05 S. Sirca BPM A/B, beam on target snapshot, run 20681
213510 11/10/07 10:13 S. Sirca Hall A General Tools
213509 11/10/07 10:12 Sparveris target screens
213506 11/10/07 09:43 S. Sirca Ozone/radiation strip charts
213505 11/10/07 09:43 S. Sirca BPM/BCM strip charts
213501 11/10/07 09:28 Sparveris arrow in target screen
213498 11/10/07 09:26 V. Sulkosky Re: BMPA Still troublesome
213493 11/10/07 09:10 Xinhu Yan Total Shower peak position after calibration
213490 11/10/07 08:56 eschulte Raster current at 276.664 MeV/c, 120 Deg.
213479 11/10/07 08:00 Z.-E. Mezian Shift Summary
213465 11/10/07 06:36 xinhu Target Dummy --->>Carbon
213462 11/10/07 06:27 xinhu Target H2-->>Dummy
213457 11/10/07 05:47 xinhu Target He-->>Pb
213454 11/10/07 05:40 Z.Meziani Momentum change on both spec.
213449 11/10/07 05:21 xinhu Target Pb-->>He4
213444 11/10/07 04:53 xinhu Target H2-->>Pb
213443 11/10/07 04:53 Z. Meziani ROC 3 disconnect
213442 11/10/07 04:42 Xinhu ROC3 is disconnected,we will reboot it.
213438 11/10/07 04:27 xinhu Target Dummy-->>H2
213433 11/10/07 04:14 xinhu Target Carbon-->>Dummy
213430 11/10/07 04:00 xinhu Right arm vdc u2 efficiency is okay
213429 11/10/07 03:53 xinhu Target Fe--->>>Carbon
213424 11/10/07 03:37 Meziani BPM/BCM charts
213423 11/10/07 03:35 Meziani Radiation-Ozone Charts
213422 11/10/07 03:32 Z. Meziani Current chart
213421 11/10/07 03:31 Meziani Hall A tools
213414 11/10/07 03:07 xinhu Target Carbon--->>Fe
213413 11/10/07 03:02 Z. Meziani VDC U2 plane efficiency
213409 11/10/07 02:24 xinhu Target Dummy--->>>Carbon
213406 11/10/07 02:08 xinhu Target He4--->>>Dummy
213401 11/10/07 01:50 xinhu Target Pb--->>>He4
213394 11/10/07 01:26 jpchen LeCroy HV ioc reboot
213389 11/10/07 01:06 xinhu target H2--->>Pb
213384 11/10/07 00:50 xinhu Target Fe--->>>H2
213381 11/10/07 00:46 eschulte BMPA Still troublesome
213380 11/10/07 00:45 xinhu Yan target carbon--->>Fe
213375 11/10/07 00:35 xinhu Yan target charts
213374 11/10/07 00:34 jpchen right spectrometer ioc rebooted
Friday
213362 11/09/07 23:59 cusanno Swing Shift Summary
213359 11/09/07 23:34 cusanno Right Arm momentum 376.66 MeV
213357 11/09/07 23:23 cusanno Left Arm momentum 464.72 MeV
213328 11/09/07 21:08 cusanno new momentum settings
213317 11/09/07 20:22 garibaldi ESR stripchart
213316 11/09/07 20:21 garibaldi target stripcharts
213314 11/09/07 20:20 garibaldi target screen
213313 11/09/07 20:15 cusanno BPM/BCP
213311 11/09/07 20:12 cusanno Compton and Radiation
213310 11/09/07 20:10 cusanno Hall A Tools
213307 11/09/07 19:50 re: cavity BPM/BCM cabling
213305 11/09/07 19:42 cusanno Right Arm Momentum 351.189 MeV
213301 11/09/07 19:25 cusanno changed Left momentum
213295 11/09/07 18:59 cusanno problem with TS on Left Arm
213292 11/09/07 18:35 cusanno end of run 2152 failed
213287 11/09/07 18:18 cusanno run 2150 and 2151 have to be summed
213283 11/09/07 18:02 cusanno New momentum settings
213281 11/09/07 17:53 eschulte Target position for Run 20629
213280 11/09/07 17:50 cusanno run 20644 is junk
213276 11/09/07 17:43 cusanno ADC 20 on Pion rejector missing
213273 11/09/07 17:39 cusanno some runs have bad target labels in automatic entries
213268 11/09/07 17:15 Ole Hansen adaql10 is back
213255 11/09/07 16:28 Sirish Compton Polarimeter
213252 11/09/07 16:04 S. Sirca Day shift SUMMARY
213248 11/09/07 15:54 brads Right Q1 current oscillation is known issue
213241 11/09/07 15:32 S. Sirca Right HRS Q1 oscillates slightly
213234 11/09/07 14:52 V. Sulkosky LHRS Gas Cerenkov PMT # 10
213229 11/09/07 14:15 S. Sirca BPM/BCM screens
213228 11/09/07 14:13 S. Sirca Ozone/radiation
213225 11/09/07 14:03 S. Sirca On rates, wrong column
213223 11/09/07 14:01 S. Sirca Question on rates
213221 11/09/07 13:46 Lingyan Zhu target screens
213194 11/09/07 12:08 V. Sulkosky BPM/Raster Fixed
213193 11/09/07 12:08 S. Sirca Hall A General Tools
213188 11/09/07 11:44 S. Sirca BPM A/B raw signals, run 2128
213179 11/09/07 10:24 Lingyan Zhu target IOC reboot
213176 11/09/07 09:00 brads Left BPM observations
213160 11/09/07 08:03 jpchen owl shift summary
213153 11/09/07 07:19 jpchen right dipole is set
213148 11/09/07 06:47 jpchen beam back
213147 11/09/07 05:58 jpchen right dipole.
213146 11/09/07 05:46 jpchen Left p set
213145 11/09/07 05:38 jpchen left p to 813.763 MeV, target top He4.
213144 11/09/07 05:20 jpchen beam off
213141 11/09/07 04:52 A. Camsonne NaI HV changed
213140 11/09/07 04:44 jpchen right dipole "quench"
213133 11/09/07 04:25 jpchen x scan on Pb target
213118 11/09/07 03:55 jpchen Left p changed to 215.105 MeV
213117 11/09/07 03:53 A. Camsonne Left arm BPM readout
213116 11/09/07 03:50 jpchen controlled access ended
213115 11/09/07 03:46 jpchen BPM/BCM strip chart
213114 11/09/07 02:27 jpchen radiation/ozone strip chart
213113 11/09/07 02:23 jpchen requested controlled access
213102 11/09/07 01:24 jpchen raster size, left has a problem
213099 11/09/07 01:07 jpchen beam x scan with empty target--> x is offset by 1 mm
213098 11/09/07 00:51 A. Camsonne Compton issue
213097 11/09/07 00:51 A. Camsonne Compton crash
213090 11/09/07 00:11 miham Target page
213089 11/09/07 00:05 miham Missing end of run 2101
213085 11/09/07 00:00 Eric Voutier Shift Summary
Thursday
213078 11/08/07 23:52 Eric Voutier Bull eyes -2 + 2
213073 11/08/07 23:47 Eric Voutier Bull Eyes (-2,-2)
213072 11/08/07 23:45 cusanno Lead target issue
213059 11/08/07 23:28 miham BCM & BPM
213058 11/08/07 23:26 miham Ozone & Radiation Monitors
213057 11/08/07 23:23 miham Hall-A Tools Page
213044 11/08/07 22:07 A. Camsonne NaI HV changed
213039 11/08/07 20:44 E. Schulte Still Tuning Beam
213032 11/08/07 19:12 V. Sulkosky HV Settings LHRS
213031 11/08/07 19:04 V. Sulkosky LHRS Gas Cerenkov
213026 11/08/07 16:45 A. Camsonne New HV NaI
213023 11/08/07 16:14 A. Camsonne NaI HV changed
213022 11/08/07 16:03 S. Sirca Day shift SUMMARY
213020 11/08/07 15:27 fansler Right Q1/D1 He read back circuit
213019 11/08/07 15:14 K.Allada Target screens
213016 11/08/07 14:49 A. Camsonne NaI HV changed
213012 11/08/07 13:05 K.Allada Target screens
213011 11/08/07 12:52 S. Sirca Ozone/radiation
213010 11/08/07 12:51 S. Sirca Hall A General Tools
213009 11/08/07 11:35 A. Camsonne Readded disk 1 to DAQ
213008 11/08/07 11:13 A. Camsonne NaI channel 263
213007 11/08/07 11:09 A. Camsonne NaI HV
213004 11/08/07 10:23 Ole Hansen all data disks are up
213003 11/08/07 09:08 S. Sirca Hall in controlled access
213002 11/08/07 08:21 S. Sirca Missing (-2,+2) point in bull's eye scans
213001 11/08/07 08:21 jpchen bull's eye scan summary
213000 11/08/07 08:09 S. Sirca Bull's Eye scan (2,-2) H03B
212999 11/08/07 08:08 S. Sirca Bull's eye scan (2,-2) H03A
212993 11/08/07 07:56 kumbartki Harp scan A
212992 11/08/07 07:51 kumbartzki Harp scan
212987 11/08/07 07:12 kumbartzki Dipol Right Arm goes down
212984 11/08/07 06:56 kumbartzki Harp scan (2,2)
212983 11/08/07 06:54 kumbartki harp scan (2,2)
212980 11/08/07 04:42 jpchen FSD problem
212979 11/08/07 03:58 jpchen Harp B results for (0,0)
212978 11/08/07 03:57 jpchen Harp A results for (0,0)
212977 11/08/07 03:49 jpchen OTR for (0,0)
212974 11/08/07 03:47 jpchen harp scan for (0,0)
212973 11/08/07 03:38 jpchen harp scan results
212970 11/08/07 03:13 jpchen Right dipole "quench", ramping down now
212969 11/08/07 03:11 jpchen Compton
212966 11/08/07 02:42 Kumbartzki Harp A scan
212965 11/08/07 02:40 kumbartzki Harp scann (0,0)
212964 11/08/07 02:30 jpchen OTR for bull's eye scan (0,0)
212963 11/08/07 02:09 jpchen DAQ deadtime ok now
212960 11/08/07 01:58 jpchen main tools screen
212959 11/08/07 01:57 jpchen right NMR and dipole
212958 11/08/07 01:15 jpchen BPM/BCM strip chart
212957 11/08/07 01:14 jpchen Radiation/Ozone strip chart
212955 11/08/07 01:04 jpchen BCM calibration
212953 11/08/07 00:34 jpchen DAQ deadtime large with very low rate!
212950 11/08/07 00:17 S. Sirca Swing shift SUMMARY
212948 11/08/07 00:05 miham BCM / BPM
212947 11/08/07 00:03 miham Ozone / Radiation
212946 11/08/07 00:02 miham Hall-A Tools Page
Wednesday
212945 11/07/07 23:43 jpchen new harp scan results
212944 11/07/07 23:39 jpchen set both spectrometers to 813.763 MeV
212943 11/07/07 23:05 jpchen harp scan results
212942 11/07/07 23:03 jpchen HV for harps are on
212941 11/07/07 20:26 E. Schulte Data Logger for 0L02 Calibration
212936 11/07/07 19:50 E. Schulte Status of Hall A
212933 11/07/07 18:42 V. Sulkosky LHRS Cerenkov HV Changes
212928 11/07/07 15:47 ewing target window
212927 11/07/07 15:40 fansler RQ1 He LVL
212926 11/07/07 15:07 Xinhu Yan the Resolution for Preshower and Shower
212925 11/07/07 13:11 A. Camsonne db_run.dat
212924 11/07/07 11:26 A. Camsonne Right arm TS0 Ok now
212923 11/07/07 11:25 V. Sulkosky Left Cerenkov Positioning
212922 11/07/07 09:50 V. Sulkosky Left Gas Cerenkov PMT # 10
212919 11/07/07 07:26 R. Michaels re: R-arm TS0 problem
Tuesday
212916 11/06/07 20:53 Xinhu Yan update pedstal value for Prl1 ,Prl2 and NaI
212915 11/06/07 20:53 Xinhu Yan update pedstal value for Prl1 ,Prl2 and NaI
212914 11/06/07 19:21 V. Sulkosky Left Gas Cerenkov Positioning
212913 11/06/07 18:36 Jack Notice for Right Dipole "Quench"
212912 11/06/07 16:44 E. Schulte Swing Shift 11/06 and Owl Shift 11/07
212910 11/06/07 15:59 hyao target chart
212908 11/06/07 15:38 S. Sirca Day shift SUMMARY
212906 11/06/07 13:45 S. Sirca General tools
212905 11/06/07 13:44 Xinhu Yan Calibration of Preshower and Shower for Right arm
212904 11/06/07 13:41 S. Sirca Ozone/radiation stripchart
212901 11/06/07 11:11 E. Schulte Updated plans for the next few days
212898 11/06/07 07:54 cisbani Owl shift summary
212897 11/06/07 07:47 cisbani Move both arms to 110.719 MeV/c
212894 11/06/07 07:35 cisbani Move back left to 698.895 MeV/c
212856 11/06/07 03:35 cisbani target windows
212854 11/06/07 03:29 cisbani Target plots
212853 11/06/07 03:28 cisbani Magnet fields stripcharts
212852 11/06/07 03:28 cisbani BCM / BPM stripchart
212851 11/06/07 03:27 cisbani Ozone stripchart
212850 11/06/07 03:26 cisbani Hall A Tools
212833 11/06/07 02:06 A. Camsonne Giving up on right arm for now
212828 11/06/07 01:45 A. Camsonne DAQ Right ARM TS stuck
Monday
212810 11/05/07 23:56 Ole Hansen Shift summary
212807 11/05/07 23:30 Ole Hansen HRSR dipole ramping up
212805 11/05/07 23:25 Ole Hansen More on HRSR dipole
212803 11/05/07 23:19 Ole Hansen HRSR dipole woes
212801 11/05/07 23:04 Ole Hansen HRSR dipole cleared
212790 11/05/07 21:52 Ole Hansen HRSR dipole ramping down
212786 11/05/07 21:42 Ole Hansen HRS-R dipole quench?
212782 11/05/07 21:19 Kumbartzki Target screens
212780 11/05/07 21:13 Ole Hansen Stripcharts
212779 11/05/07 21:11 Ole Hansen Hall A Main Screen (with picture)
212778 11/05/07 21:11 Ole Hansen Hall A Main Screen
212770 11/05/07 20:26 V. Sulkosky Pedestals Updated for LHRS S1 and S2
212747 11/05/07 18:52 E. Schulte Warming the ESR heat exchanger and the energy change
212742 11/05/07 18:31 E. Schulte Lead Target Raster Current Density Issue
212711 11/05/07 16:04 S. Sirca Day shift SUMMARY
212709 11/05/07 15:59 S. Sirca General tools screens
212708 11/05/07 15:36 hyao target chart
212707 11/05/07 15:35 hyao target chart
212706 11/05/07 15:32 M. Mihovilov Ozone, radiation strip chart
212705 11/05/07 14:59 Xinhu Yan Calibration of Cerenkov for left and right arm
212704 11/05/07 14:59 Xinhu Yan Calibration of Cerenkov for left and right arm
212703 11/05/07 14:59 Xinhu Yan Calibration of Cerenkov for left and right arm
212702 11/05/07 14:36 S. Sirca Beam off for RF separator tune
212694 11/05/07 14:19 Xinhu Yan update the DB file ->"db.L.cer.dat"
212686 11/05/07 13:51 S. Sirca Left HRS DAQ question
212682 11/05/07 13:27 S. Sirca Data file name mixups
212658 11/05/07 11:56 E. Schulte Loss He Expected
212657 11/05/07 11:37 S. Sirca Calorimeter HV trip
212656 11/05/07 11:37 hyao target moved to He4 Loop1
212653 11/05/07 11:16 S. Sirca Report on cryotarget recovery
212650 11/05/07 11:08 hyao target moved from Fe to C
212645 11/05/07 10:43 A. Camsonne pedestal restored
212639 11/05/07 10:34 A. Camsonne Pedestal subtraction off
212635 11/05/07 09:27 hyao Controlled Access-target problem
212632 11/05/07 09:11 S. Sirca LHe vented, controlled access
212622 11/05/07 08:03 jpchen Owl shift summary (continue)
212621 11/05/07 08:02 jpchen Change right P to 698.895 MeV
212612 11/05/07 07:37 jpchen radiation/ozone strip chart
212611 11/05/07 07:36 jpchen main tools screen
212610 11/05/07 07:35 jpchen BPM/BCM strip chart
212598 11/05/07 06:45 jpchen right dipole ready
212590 11/05/07 06:05 jpchen Left arm ready
212589 11/05/07 06:03 jpchen owl shift summary
212588 11/05/07 06:01 jpchen Move both to 803.966 MeV
212562 11/05/07 04:47 jpchen right dipole set for 924.834 MeV
212559 11/05/07 04:32 jpchen beam general screen
212558 11/05/07 04:31 jpchen OTR picture
212552 11/05/07 03:55 zhangyi target chart
212550 11/05/07 03:50 jpchen ORT strip charts
212549 11/05/07 03:48 jpchen Radiation/ozone stripe charts
212547 11/05/07 03:46 jpchen BPM stripe charts
212546 11/05/07 03:37 jpchen right spectrometer moved to 60 degree
212543 11/05/07 03:32 jpchen left p to 924.834 MeV
212539 11/05/07 03:03 jpchen started to change right to 924.834 MeV, 60 degrees
212535 11/05/07 02:50 jpchen RF seperator problem for 5th pass
212522 11/05/07 02:25 jpchen RF portion of FFB turned off at 2:08
212519 11/05/07 02:18 jpchen re-take data for runs 1805-1814.
212516 11/05/07 02:12 jpchen left spectrometer to 60 degree
212512 11/05/07 02:08 jpchen right dipole set for p=110.719 MeV
212510 11/05/07 01:41 zhangyi NaI Detectors
212503 11/05/07 00:44 jpchen right dipole quenched
212497 11/05/07 00:00 Kumbartzki Swing Shift Summary
Sunday
212472 11/04/07 22:14 Kumbartzki replay wows
212461 11/04/07 21:52 mcnulty target screens
212444 11/04/07 20:54 Kumbartzki BPM charts
212429 11/04/07 20:14 A. Camsonne moved replay to /adaqfs/scratch/adaq/E05110
212422 11/04/07 19:27 A. Camsonne adaql10 down again moved replay to adaql6 disk
212421 11/04/07 19:20 A. Camsonne adaql10 was off
212373 11/04/07 16:00 cisbani Shift Summary
212364 11/04/07 15:50 cusanno run 20530 has extra-events
212357 11/04/07 15:06 cisbani BPM slightly out of range at the beginning of run
212344 11/04/07 13:50 cusanno end of run 1904 failed
212336 11/04/07 13:18 K.Allada target screens
212310 11/04/07 12:18 cisbani Right Momentum at 246.575 MeV/c
212306 11/04/07 11:43 xin target status
212301 11/04/07 11:31 cisbani BPM / BCM stripchart
212300 11/04/07 11:30 cisbani Ozone stripchart
212299 11/04/07 11:28 cisbani Hall A Tools
212297 11/04/07 11:21 folts tech /Dipole
212295 11/04/07 11:14 cisbani BPM jump during the run
212293 11/04/07 11:04 cisbani Right Arm Fixed
212285 11/04/07 10:29 cisbani Right Arm dipole quenched
212282 11/04/07 10:14 eschulte BPM Strip Chart
212270 11/04/07 09:10 cisbani OTR graphs
212268 11/04/07 09:01 eschulte BPM's at start of Day Shift
212264 11/04/07 08:52 cusanno wrong start-of-run entry
212244 11/04/07 07:47 cisbani Target in dummy 10 cm
212240 11/04/07 07:27 Seonho Choi Nov. 4 OWL shift summary
212239 11/04/07 07:26 R. Michaels TACC vs event length.
212230 11/04/07 07:05 R. Michaels wrong header
212217 11/04/07 06:25 Seonho Choi Beam position stability
212211 11/04/07 05:59 R. Michaels fluctations in event length
212199 11/04/07 05:19 R. Michaels more on DT glitches.
212192 11/04/07 04:59 R. Michaels deadtime glitches
212184 11/04/07 04:38 R. Michaels spot_R
212183 11/04/07 04:37 R. Michaels spot_L
212172 11/04/07 04:09 Seonho Choi Ozone/Radiation strip chart
212171 11/04/07 04:08 Seonho Choi BPM/BCM strip charts
212170 11/04/07 04:07 Seonho Choi Hall A Tools Page
212154 11/04/07 02:53 brads Stripchart settings with useful bounds
212151 11/04/07 02:47 R. Michaels scaler checks
212150 11/04/07 02:45 Seonho Choi Beam position stability at the start of beam delivery
212143 11/04/07 02:29 A. Camsonne Pedestal updated for NaI
212124 11/04/07 01:42 Seonho Choi Beam position at the start of the beam delivery
212093 11/04/07 01:01 R. Michaels correction to start-run entries
212091 11/04/07 00:49 Seonho Choi Beam position unstable at the start of beam delivery to the hall
Saturday
212076 11/03/07 23:58 Kumbartzki Swing shift summary
212071 11/03/07 23:42 jpchen Spikes in Gac C sum
212064 11/03/07 23:17 jpchen VDC problems for runs 1802-1814
212059 11/03/07 22:58 jpchen beam bad for run 1822
212056 11/03/07 22:37 jpchen Left Gas Cherenkov ch 10 and spikes in sum
212053 11/03/07 22:33 jpchen Bad VDC spectra was due to noisy NaI
212050 11/03/07 22:26 K.Allada run 1822 and 20483 are junk
212043 11/03/07 22:16 Kumbartzki more to beam position change
212040 11/03/07 22:08 jpchen Beam quality
212039 11/03/07 22:07 jpchen BMPA drifting
212038 11/03/07 22:06 A.Camsonne Run 1806 noisy
212037 11/03/07 22:04 Kumbartzki Beam positions changed
212036 11/03/07 22:04 jpchen BPMA also centered to (0,0).
212034 11/03/07 22:02 jpchen/camso Compton rate high, HV off
212029 11/03/07 21:56 jpchen Target OTR out
212025 11/03/07 21:46 K.Allada run 1818 target is 4He
212012 11/03/07 20:39 A. Camsonne NaI HV Channel 53 off
211992 11/03/07 19:30 eschulte HRS-L VDC looks bad
211988 11/03/07 19:21 brads OTR looks saturated => OTR data are useless
211979 11/03/07 18:39 eschulte BCM & BPM for Swing Shift
211978 11/03/07 18:38 eschulte ozone & radiation monitor strip chart for swing shift
211977 11/03/07 18:38 Kumbartzki Target charts
211976 11/03/07 18:37 eschulte compton & otr for swing shift
211975 11/03/07 18:36 eschulte Hall A tools for Swing Shift
211969 11/03/07 18:19 eschulte OTR @ 50 uA with raster on
211967 11/03/07 18:18 E. Schulte Back to taking data after along interruption of service
211966 11/03/07 18:16 eschulte New raster spot scans!
211957 11/03/07 18:03 eschulte Raster after beam profile retuning
211951 11/03/07 17:36 Kumbartzki New OTR and Harp
211950 11/03/07 17:14 Kumbartzki OTR
211949 11/03/07 16:16 A. Camsonne NaI HV changed
211946 11/03/07 16:05 cisbani Shift Summary
211945 11/03/07 16:00 E. Schulte OTR Out during these runs
211944 11/03/07 15:58 eschulte Runs before Day Shift Harp Scan
211937 11/03/07 15:24 cisbani General Accelerator Parameters
211934 11/03/07 15:04 cisbani Raster on dummy 15
211926 11/03/07 12:35 cisbani adaql2:/data1 is now out of the DAQ
211922 11/03/07 12:16 xqian Target Status
211921 11/03/07 12:13 cisbani BCM and BPM stripchart
211920 11/03/07 12:11 cisbani Ozone stripchart
211919 11/03/07 12:09 cisbani Hall A Tools
211918 11/03/07 11:08 cisbani adaql2:/data1: one disk faulty in redundant array
211917 11/03/07 10:50 cusanno Beam expected around noon
211916 11/03/07 10:01 cusanno machine hacsbc2 reboot
211915 11/03/07 09:58 cisbani Beam delivery delayed
211914 11/03/07 09:39 cisbani No harp scan, MCC has problem
211913 11/03/07 09:37 cisbani Hacsbc2 rebooted by himself
211911 11/03/07 08:05 Seonho Choi Nov. 3 OWL shift summary (right arm)
211910 11/03/07 08:03 Seonho Choi Nov. 3 Owl shift summary (Left Arm)
211909 11/03/07 07:17 mcnulty target charts
211908 11/03/07 06:12 Seonho Choi Beam taken away for Hall B separation
211901 11/03/07 05:57 Seonho Choi spot_R after beam back
211900 11/03/07 05:55 Seonho Choi spot_L after beam back
211897 11/03/07 05:40 Seonho Choi Harp scan after beam is back (IHA1H03B)
211896 11/03/07 05:38 Seonho Choi Harp Scan after beam back (IHA1H03)
211895 11/03/07 05:24 Seonho Choi Beam will be back earlier than expected
211894 11/03/07 05:21 Seonho Choi BPM and BCM stripcharts
211893 11/03/07 05:20 Seonho Choi Compton and OTR Stripchart
211892 11/03/07 05:20 Seonho Choi Ozone and radiation strip chart
211891 11/03/07 04:40 Seonho Choi Beamstudies
211890 11/03/07 04:27 Seonho Choi Nov. 3 Owl mid-term summary (Right Spectrometer)
211889 11/03/07 04:25 Seonho Choi Nov. 3 Owl shift mid-term summary (Left spectrometer)
211868 11/03/07 03:09 Seonho Choi Hall A General Tools Screenshot
211832 11/03/07 00:53 Seonho Choi Compton Detector HV off
211823 11/03/07 00:05 Solvignon Shift summary
Friday
211818 11/02/07 23:53 Solvignon Left cerenkov PMT 10 too low
211815 11/02/07 23:47 eschulte Compton Rates
211812 11/02/07 23:41 eschulte OTR screen after pass change
211808 11/02/07 23:31 jpchen Target motion to loop2 disabled for now
211806 11/02/07 23:25 E. Schulte Possible Monticello Problems
211799 11/02/07 23:11 Solvignon Harp scan result for HA1H03A
211798 11/02/07 23:10 eschulte Harp Scan Results
211797 11/02/07 23:10 Solvignon Harp scan result from HA1H03B
211794 11/02/07 23:04 Solvignon Spot++
211789 11/02/07 22:44 Kumbartzki more target screens
211788 11/02/07 22:39 Kumbartzki Target status
211787 11/02/07 22:38 Solvignon BPM and BCM stripcharts
211786 11/02/07 22:38 Solvignon Compton and OTR stripchart
211785 11/02/07 22:36 Solvignon Ozone and radiation stripchart
211784 11/02/07 22:36 Solvignon Hall A tools
211783 11/02/07 22:21 Solvignon Incorrectly label runs
211782 11/02/07 21:55 Solvignon Raster setup doesn't make sense
211781 11/02/07 21:53 eschulte rate in comption
211780 11/02/07 21:39 Solvignon Harp scan problem for B
211769 11/02/07 21:14 Solvignon Raster size with setting of 1 pass
211760 11/02/07 20:35 Solvignon Tuned beam
211759 11/02/07 19:20 V. Sulkosky Gas Cerenkov (LHRS) HV
211758 11/02/07 19:14 V. Sulkosky HV Changes for RHRS
211756 11/02/07 18:21 Solvignon iocha22 is back up
211755 11/02/07 18:16 Solvignon Beam status
211752 11/02/07 17:59 Solvignon iocha22 not responding
211751 11/02/07 17:56 brads 1H03B Harp encoder adjustment
211748 11/02/07 17:13 A. Camsonne NaI work
211745 11/02/07 16:44 brads 1458/1461N HV trip problem
211744 11/02/07 16:13 E. Schulte Beam Current on Lead Target
211743 11/02/07 16:03 averett day shift summary con't
211738 11/02/07 10:51 brads Harp 1H03B problem actually stuck bit in the encoder(?) + Caveat for Harp analyser tool
211737 11/02/07 10:43 brads Paranoid question re: old Harp scan software and hardware adjustments
211728 11/02/07 09:51 E. Schulte Bad NaI channels as of Day Shift 11/02/2007
211725 11/02/07 09:42 gilman yellow alarm transient
211722 11/02/07 09:41 E. Schulte Rebooted & Restarted Compton CODA
211717 11/02/07 09:24 averett Beg. Day Shift Summary
211710 11/02/07 09:06 E. Schulte Harp Scan Data Analyzer
211680 11/02/07 08:03 K.Allada owl shift summary
211671 11/02/07 07:42 xinhu Target Pb-->>H2
211666 11/02/07 07:33 K.Allada OTR screen
211663 11/02/07 07:24 xinhu target He--->>Pb
211654 11/02/07 06:55 Xinhu Target Fe--->>>He
211649 11/02/07 06:45 xinhu Target Carbon-->>Fe
211646 11/02/07 06:32 Xinhu Target Dummy-->>Carbon
211643 11/02/07 06:21 Xinhu target He--->>Dummy
211634 11/02/07 05:25 Xinhu target Pb--->>He
211631 11/02/07 05:12 Xinhu Target 2H--->>Pb
211628 11/02/07 05:07 Xinhu target moved to 2H
211619 11/02/07 04:42 Xinhu target moved to 12C
211614 11/02/07 04:33 Xinhu Target He-->>Fe
211607 11/02/07 04:06 Xinhu Target --->>He
211601 11/02/07 03:50 Xinhu Target-->>H2
211598 11/02/07 03:36 Xinhu Target status
211590 11/02/07 02:55 Xinhu Target -->>Fe
211587 11/02/07 02:40 Xinhu Target He--->>C12
211586 11/02/07 02:39 jpchen energy lock switched to Hall A
211580 11/02/07 02:20 jpchen main tools screens
211577 11/02/07 02:19 jpchen OTR Compton strip charts
211576 11/02/07 02:17 jpchen BPM/BCM strip charts
211575 11/02/07 02:16 jpchen radiation/ozone strip charts
211570 11/02/07 02:06 Xinhu Target -->>H2
211565 11/02/07 01:51 Xinhu Target -->>Dummy
211555 11/02/07 01:20 Xinhu Yan Target -->>12C
211552 11/02/07 01:07 Xinhu Yan Target charts
211551 11/02/07 01:07 xinhu/jpchen problem channels on NaI
211550 11/02/07 01:06 Xinhu Yan Target moved to Dummy
211547 11/02/07 00:53 Xinhu Yan Target moved from Pb to H2
211544 11/02/07 00:38 jpchen Compton restarted
211542 11/02/07 00:10 cisbani Swing shift summary
Thursday
211536 11/01/07 23:52 chiranjib TARGET---->>He
211535 11/01/07 23:45 cisbani Left momentum 206.972 MeV/c
211530 11/01/07 23:32 Chiranjib TARGET------>>>> Fe
211525 11/01/07 23:23 Chiranjib TARGET---->>C12
211508 11/01/07 23:04 A. Camsonne Not very useful
211507 11/01/07 23:03 Chiranjib TARGET---->> DUMMY
211498 11/01/07 22:51 chiranjib TARGET---->>H2
211493 11/01/07 22:40 chiranjib TARGET---->>Pb
211490 11/01/07 22:21 Chiranjib TARGET------>> He
211487 11/01/07 22:03 Chiranjib TARGET ------>> Fe
211486 11/01/07 22:00 cisbani Righ momentum 252.838 MeV/c
211483 11/01/07 21:53 Chiranjib TARGET----->> C12
211480 11/01/07 21:43 Chiranjib TARGET------->>> DUMMY
211473 11/01/07 21:29 Chiranjib TARGET--->>>H2
211460 11/01/07 21:16 Chiranjib TARGET----->> Pb
211455 11/01/07 20:57 Chiranjib TARGET---->>> He
211454 11/01/07 20:51 cisbani Left momentum 279.453 MeV/c
211449 11/01/07 20:39 Chiranjib TARGET------>>> Fe
211444 11/01/07 20:19 Chiranjib TARGET------>>> C12
211433 11/01/07 20:04 chiranjib TARGET------>>> DUMMY
211432 11/01/07 20:01 Adaq Raster f or He run, sort of halo
211423 11/01/07 19:51 Chiranjib TARGET SCREENS II
211422 11/01/07 19:49 Chiranjib TARGET----->> H2
211421 11/01/07 19:48 cisbani Right Arm Momentum at 341.381 MeV/c
211420 11/01/07 19:46 Chiranjib TARGET SCREENS I
211419 11/01/07 19:46 cisbani Left Momentum at 308.869 MeV/c
211414 11/01/07 19:39 cisbani BCM / BPM stripcharts
211413 11/01/07 19:38 cisbani Ozone stripchart
211412 11/01/07 19:37 Chiranjib TARGET------>>>> Pb
211401 11/01/07 19:17 Chiranjib TARGET---->>> He
211398 11/01/07 18:49 A. Camsonne NaI status
211397 11/01/07 18:46 Chiranjib TARGET------>>>> Fe
211396 11/01/07 18:44 cisbani Hall A Tools
211391 11/01/07 18:36 Chiranjib TARGET------->>>> C12
211386 11/01/07 18:18 Chiranjib TARGET------>>>> DUMMY
211381 11/01/07 18:07 Chiranjib TARGET----->> Pb
211368 11/01/07 17:39 cisbani left momentum 377.316 MeV/c
211363 11/01/07 17:23 cisbani raster on Fe run
211362 11/01/07 17:23 Chiranjib TARGET------>>>> He
211361 11/01/07 17:16 jpchen pre-scale T2 (T4) before pre-scale T1 (T3)
211358 11/01/07 17:07 cisbani Right momentum at 417.033 MeV/c
211355 11/01/07 17:03 Chiranjib TARGET----->>>> Fe
211350 11/01/07 16:50 Chiranjib TARGET----->>> C12
211348 11/01/07 16:42 Chiranjib TARGET------>>>> DUMMY
211342 11/01/07 16:18 cisbani Left Momentum 460.932
211341 11/01/07 16:16 Chiranjib TARGET ----->>> Pb
211331 11/01/07 15:57 gilman day shift summary
211330 11/01/07 15:54 hyao target moved to He-loop1-10cm
211327 11/01/07 15:49 eschulte Pb target test, double raster size
211316 11/01/07 15:34 eschulte Let's Try this again: Additional Apparently not-working NaI channels
211315 11/01/07 15:33 hyao target moved to Loop3-Pb-15cm
211310 11/01/07 15:22 hyao target moved to dummy-10cm
211307 11/01/07 15:20 eschulte Additonal Apparently Non-functioning NaI channels
211306 11/01/07 15:19 gilman shift checklist screens - bpm and bcm
211303 11/01/07 15:13 hyao target moved to Beam right C
211302 11/01/07 15:13 gilman ozone for shift checklist
211299 11/01/07 15:11 hyao target chart 2
211298 11/01/07 15:10 hyao target charts 1
211297 11/01/07 15:08 gilman halla tools
211290 11/01/07 15:04 gilman harp 1h03b
211289 11/01/07 14:59 gilman harp scan
211288 11/01/07 14:48 J.Segal /A. HV 1458 limit
211287 11/01/07 14:45 gilman beam back - spot looks good
211281 11/01/07 14:13 brads Rebooted RHRS magnet controls ioc (iocha16)
211280 11/01/07 13:57 A. Camsonne NaI channels on Lecroy 1458
211279 11/01/07 13:19 A. Camsonne Channel 37 on patch left arm
211274 11/01/07 12:23 A. Camsonne NaI HV increased
211273 11/01/07 11:59 brads Right NMR readback fixed
211271 11/01/07 11:47 Y. Oh/ A. Ca HV NaI
211270 11/01/07 11:10 brads Modified runstart.tcl scripts
211253 11/01/07 10:04 gilman 120 degree setting done
211246 11/01/07 09:49 hyao target moved to beam right Fe
211241 11/01/07 09:41 hyao target moved to Beam right C
211236 11/01/07 09:32 hyao target moved to dummy-10cm
211233 11/01/07 09:25 eschulte bpm strip chart for late owl shift
211230 11/01/07 09:22 hyao target moved to loop3-Pb-15cm
211221 11/01/07 08:50 hyao target moved to H3-Loop1-10cm
211216 11/01/07 08:38 hyao target moved to Loop3-Pb-15cm
211211 11/01/07 08:26 hyao target moved to H2-Loop3-15cm
211200 11/01/07 08:02 jpchen Owl shift summary
211196 11/01/07 07:57 jpchen right dipole almost ready for P=116.986 MeV
211195 11/01/07 07:41 jpchen Start to change left P to 474.672 MeV.
211194 11/01/07 07:37 zhangyi raster tail
211188 11/01/07 07:11 jpchen target otr in again
211184 11/01/07 06:56 jpchen start to change right P to 116.986 MeV
211172 11/01/07 06:27 jpchen Left P set to 109.074 MeV
211164 11/01/07 06:08 zhangyi raster tail
211157 11/01/07 05:54 jpchen right dipole set for 134.573 MeV
211151 11/01/07 05:28 zhangyi raster tail
211149 11/01/07 05:24 jpchen change left p to 125.472 MeV
211142 11/01/07 04:46 jpchen start to change right P to 134.573 MeV
211129 11/01/07 04:09 jpchen Left P change to 144.335 MeV
211118 11/01/07 03:40 jpchen tools screen
211115 11/01/07 03:34 jpchen right dipole set for P=166.034 MeV
211114 11/01/07 03:33 jpchen/yizha raster, beam tail in y?
211107 11/01/07 02:59 jpchen Change Left P to 154.805 MeV
211099 11/01/07 02:35 jpchen OTR
211097 11/01/07 02:26 jpchen start to change right P to 166.034 MeV
211094 11/01/07 02:25 jpchen target theta tail
211087 11/01/07 02:02 jpchen target screens
211082 11/01/07 01:52 jpchen left P to 178.078 MeV
211081 11/01/07 01:48 jpchen Compton
211078 11/01/07 01:41 jpchen BPM and BCM
211077 11/01/07 01:40 jpchen radiation and ozone
211076 11/01/07 01:37 jpchen Hall A tools screen
211065 11/01/07 01:11 jpchen changed right p to 204.851 MeV
211062 11/01/07 01:00 jpchen raster for run 1595
211059 11/01/07 00:38 V. Sulkosky Plots for Run # 1497
211056 11/01/07 00:26 V. Sulkosky Cuts Placed on Figure 2 (bottom)
211054 11/01/07 00:18 V. Sulkosky Data with Bad Beam Tune (509 MeV/c)
211052 11/01/07 00:14 A. Camsonne HV channel 136 was low
211051 11/01/07 00:12 jpchen Left p changed to 190.996 MeV
211050 11/01/07 00:07 Chiranjib TAREGT ------>>> He
211048 11/01/07 00:03 cisbani Swing Shift Summary